# A new economic and financial theory of money

### v5

@@ VII Solution of the Hamilton-Jacobi-Bellman equation
is quite interesting. Optimal control as done in AI with Deep Q-Learning (DQN) (Mnih et al., 2015) is based on parametric estimations of a Q-function
@@
Given this theory, we move on to the practical application of it to control the system. This application will use the concepts of Artificial Intelligence (Hastie et al., 2009; Sugiyama, 2015; Goodfellow et al., 2016), as interpreted by Glinsky and Sievert (2024). First, construct a dataset by either doing an ensemble of simulations of the system or by observing the system. It will be assumed that the system has a small number of dimensions. Most of the systems of interest present themselves in the domain of collective fields $f(x)$ that are elements of a Hilbert space not $q(\tau)$ with a small number of components. How to de-convolute from the domain of the collective motion, the field $f(x)$, to the domain of the individual, $q(\tau)$, (that is from a Hilbert space to $\mathbb{C}^{n}$) using the Heisenberg Scattering Transformation (HST) and a Principal Components Analysis (PCA) will be discussed at the end of this section. This transformation can be done because the collective acts as one, because of the correlation or synchronization specified by the S-matrix $S_{m}$ given in Eq.  91, which are the derivatives of the analytic Hamiltonian $H(\beta)$ of the individual. It is helpful to apply an external force $F_{\text{ext}}(p,q)$ to the system being simulated or observed to sample phase space more efficiently. A good choice would be a dissipation or a random diffusion which samples phase space well, as the system gradually relaxes to the stable equilibriums. It is also good to apply an external force that is constructed to keep the dynamic trajectory in the vicinity of the unstable local maximums, that is stabilizes the unstable equilibriums. The set of variables that should be recorded are $\tau$, in addition to variables that are related to the state $q(\tau)$ and the co-state $p(\tau)$.
@@
Given the dataset, train a neural network with an architecture that matches the structure of the solution to the HJB equation to estimate: (1) the encoding of the $p$ and $q$ coordinates into the $P(p,q)$ and $Q(p,q)$ coordinates that are the solution to the HJB equation as well as the decoding of $P$ and $Q$ to $p(P,Q)$ and $q(P,Q)$, (2) the value function $W(q,P)$ that generates these canonical transformations and is related to the imaginary part of the analytic Hamiltonian as shown in Eq. 83, (3) the mapping of $P$ to the real part of the analytic Hamiltonian $E_{P}(P)$, (4) the frequency $\omega_{Q}(P)$, (5) the policy $\pi(q,P)$ and (6) the analytic advance of $P$ and $Q$ given in Eqns. 79 and 78. Multi Layer Perceptrons (MLPs) (Goodfellow et al., 2016) are used to approximate some of the functions. The derivative functions are calculated by back propagating the MLPs. This architecture is shown in Fig. 9. It is important that Rectified Linear Units (ReLUs) are used as activation functions in the MLPs because the MLPs are approximating analytic functions which are maximally flat but do have a limited number of singularities where the derivative is discontinuous. MLPs with ReLUs are very good at doing this since they are universal piecewise linear approximators with discontinuities in the derivative.
@@
Figure 9: The neural network architecture to estimate the solution of the Hamilton-Jacobi-Bellman equation. It starts with a decoder from $(p,q)$ to $(P,Q)$ generated by the value function $W(q,P)$ related to the imaginary part of the analytic function $H(\beta)$ by Eq. 83, a mapping to the real part of the analytic function $H(\beta)$ given by $E_{P}(P)$, the frequency $\omega_{Q}(P)$, the policy $\pi(q,P)$, the analytic mapping to advance $(P,Q)$, and a decoder from $(P,Q)$ to $(p,q)$. The universal function approximator is a Multi Layer Perceptron (MLP).
@@
Given the function $E_{P}(P)$ estimated in the workflow shown in Fig. 9, $P^{*}$ can be found with a high performance root finder, both stable and unstable equilibriums. The equilibrium policy can then be estimated as $\pi^{*}(q)=\pi(q,P^{*})$ and the equilibrium value as $V^{*}(q)=W(q,P^{*})$. One now has estimated the $\beta^{*}$ which are viewed different ways by different technical disciplines. These are: (1) the ground states of quantum field theory (Weinberg, 2005), (2) the attractive manifolds of nonlinear dynamics (Lichtenberg and Lieberman, 2010), (3) the emergent behaviours and self organizations of complex systems (Gros, 2015), (4) the Taylor relaxed states (Taylor, 1986) and BGK modes (Bernstein et al., 1957) of plasma physics, (5) the poles and branch cuts of control theory and complex analysis (Nehari, 2012), and fundamentally (6) the homology classes of the topology of the dynamic manifold or the geometry of the physics (Frankel, 2012). The equilibrium values $\beta^{*}$ need not be points. They can be algebraic structures, if $n>1$.
@@
There are also unstable equilibriums where $\omega^{\prime}_{Q}<0$, that is local maximums. These unstable equilibriums are also known as saddle points or x-points or strange attractors that first attract the trajectory to them with the system staying near them for a long time since $\omega_{Q}=0$ (a meta stable state). But, eventually the system will start to move away from the strange attractor, being repelled away from it. The trajectory will eventually approach and be attracted back to the strange attractor or might approach and be attracted to another strange attractor (or several other strange attractors) before being repelled by the last strange attractor and returning to the first strange attractor. An example of a set of two strange attractors could be a bull and a bear market. These saddle points will need to be stabilized by application of an external conservative force, then cooled by a matched non-conservative thermal force, otherwise the trajectory will evolve into the basin of attraction of a stable attractor and eventually relax to that attractor. While small amounts of energy are needed to move from strange attractor to strange attractor, a significant amount of energy is needed to move from one attractor to another attractor or strange attractor. One needs to climb out of the basin of attraction or potential well, that is climb back up to the mountain pass (saddle point) out of the basin before descending into another basin. When these strange attractors are stabilized, there is still a local minimization of the economic action to efficiently accomplish the task, but the sub-economy is kept at a global sustainable maximum.
@@
Figure 14: Graph showing the energies $E(P^{*})$, where $\omega_{Q}(P^{*})=0$, as a function of the regularization or dissipation $\nu$ and the strength of the ponderomotive stabilization force $f_{0}$. The thick green line shows the unstable equilibrium or x-point or local maximum at infinity with energy $E_{\text{x}\infty}$. The thick red line shows the stable equilibrium or o-point or Nash Equilibrium at zero with energy $E_{\text{o0}}$. This is the point of severe economic depression. The thin green line shows the local maximum at $P^{*}_{\text{x}}$ with energy $E_{\text{x}}$. The thin red line shows the Nash Equilibrium at $P^{*}_{\text{o}}$ with energy $E_{\text{o}}$. This is the point of economic depression. The thin dashed green lines show the local maximums of the stabilization force with energies $E_{\text{+}}$ and $E_{-}$. The thin dashed red line shows the Nash Equilibrium of the stabilization force with energy $E_{\text{s}}$. This is the point of economic prosperity. The upper limit on viscosity so that the singularities $P^{*}_{\text{x}}$ and $P^{*}_{\text{o}}$ are not destroyed is $\nu<\nu_{\text{cr}}$.
@@
Now to the postponed, but important issue, that the system does not present itself in the low finite dimensional dynamical state $q$ and canonically conjugate momentum $p$ or co-state of the individual, but as an infinite dimensional field $f(x)$ and its canonically conjugate field momentum $\pi(x)$ of the collective, where $x$ is the base manifold or space with metric $\tau(x)$. Another way of looking at this is that nature presents itself as a convoluted form of the dynamical variables that needs to be de-convoluted or decoded into the dynamical variables. It is very common that $x=t$ and $\tau(t)=t$, but could be as sophisticated as $x=(x,y,z,t)$ and $\tau(x,y,z,t)=\sqrt{t^{2}-(x^{2}+y^{2}+z^{2})/c^{2}}$ (that is relativistic dynamics on space-time). The transformation from the infinite dimensional Hilbert space with coordinates $(\pi,f)$ to the finite dimensional $\mathbb{C}^{n}$ space with coordinates $(p,q)$, has been an ongoing challenge to physics and complex system analysis. It has manifested itself as the renormalization challenge to physicists that was first addressed by Ken Wilson (Wilson, 1971) and has been the subject of two Nobel prizes. There have been recent developments on this subject that have culminated in the Heisenberg Scattering Transformation (HST). The HST is a closed form, specified Convolutional Neural Network (CNN) or Wavelet Conditional Renormalization Group (WC-RG) (Marchand et al., 2022) which is a transformation from canonical $(\pi,f)$-field phase space to $\mathbb{C}^{n}$ with coordinates $\beta$. When combined with the specification of the analytic function $H(\beta)$ or the equilibrium points (topological homology classes) $\beta^{*}$, it is a Generative Pretrained Transformer (GPT) (Li et al., 2022), and when further deployed as a controller it is a Deep Reinforcement Learning (DRL) (Bertsekas and Tsitsiklis, 1996; Sutton and Barto, 2018; Yoon et al., 2021; Wang et al., 2021; Mnih et al., 2015). The HST will be used as a pre-processor, and the inverse HST (iHST) will be used as a post-processor to the previously discussed analysis, as shown in Fig. 16. A detailed discussion of the relationship of our optimal control methodology to DRL can be found in Sec. X, where particular attention is paid to practical differences between our optimal control methodology and DRL.
@@
Details of the HST can be found in Glinsky and Maupin (Glinsky and Maupin, 2023; Glinsky, 2023a), along with an example of how it can be combined with an MLP and PCA to form a surrogate model for resistive MagnetoHydroDynamics. The equation for the HST is
@@
What the control method of this section does is move the equilibriums from their uncontrolled natural positions to the desired positions according to the reward function $R(q)$, not $R_{0}(q)$, and stabilizes and cools a more desirable (with more economic activity) saddle point or strange attractor.
We digress to note the present use of the prime interest rate $\nu_{\text{prime}}$ in economic control. One of the ways that the money supply is traditionally controlled is by a central reserve bank loaning money at a set rate. If this rate is lower than the natural prime interest rate, the money supply will increase as money is borrowed from the central reserve bank. This is what is called quantitative easing and puts inflationary pressure on the economy. The money supply will decrease if the set rate is greater than the prime rate. This reduction in money supply will counter inflation. As will be discussed in Sec. XV, this is a poor control system since the loaning of money causes a long-term consistent reduction in the money supply due to the coupon payments, countering the original increase in the money supply.
@@ VIII FASER: the LASER of financial systems
The financial FASER works the same way. The financial sub-system needs to be in a meta stable excited state, enjoying the irrational exuberance of a bull market. This meta stable state is likely an unstable saddle point or strange attractor that takes an infinite amount of time to approach since $\omega_{Q}^{*}=0$. The market is then pumped into an even more excited bull state that is so over priced that it triggers a bear market and the market falls back into its ground (that is) stable state. The financial gain (or profit/loss) is $E_{\text{gain}}$ which is positive for $\Delta E_{p}<\Delta E_{d}$. If the system is already in the ground state $E_{\text{gain}}=-\Delta E_{p}<0$ – a guaranteed money losing proposition pumping energy from the entity doing the “pump and dump” to the financial sub-system.
@@ X Relationship to Deep Reinforcement Learning
Once these surrogates are formed, the equilibrium points $P^{*}$ can be efficiently found in 10’s of seconds. Not only are the stable equilibriums (that is local minimums, o-points, or attractors) found, but the more desirable unstable equilibriums (that is local maximums, x-points, saddle points, or strange attractors) are also found. These unstable equilibriums can then be stabilized and protected from external forces by the use of Eq. 100 or Eq. 99. The reward function can also be modified as shown in Eq. 88. DRL uses a much slower gradient decent that needs to calculate two time-consuming minimums or maximums for each sample.
@@ XI Quantum field nature of financial systems
The quantum field structure of financial systems comes about from real-time or space-like measurement limitations to those of a statistical nature, so that the Born Rule is applicable. The Born Rule states that if a statistical measurement is made on a system, the measurement must exert a force on the system that will change the system dynamics – the measurement is destructive. Only statistical, that is destructive measurements, can be made on relativistic systems, systems that present themselves in the convoluted field domain, or systems where non-destructive observations are not available. Practically, this comes about from the time that it takes to communicate, sum, or assimilate the measurements. For instance, financial statements are only issued quarterly or annually, in a retrospective nature. Even if they are reported real-time, they are still only retrospective. The destructive measurement forces are most often non-conservative viscous forces, but can be conservative forces that erect a potential barrier that redirect, not extract energy from, the dynamical trajectory.
@@
The current banking system suffers from an inability to make detailed non-destructive observation of the system. It results in many different money supplies $M$, that are all partial or indirect estimates of the true money supply, at best. The true money supply must be destructively measured as discussed earlier in this section. It is aggravated by the liquidity uncertainty generated by the banks not holding most of the balance of its customers savings accounts in cash. This requires that the depositors treat their deposits statistically, and make statistical measurements. The common statistical measurement of the depositors is requiring the bank to make regular interest payments. In more extreme cases, it can result in the depositors making a very destructive measurement, a bank run where the depositors require the bank to return all their deposits, that results in bank failures, that is bank death, as discussed in Sec. XV. In both cases, dissipation is added to the economy, degrading the economy’s geometry and performance, and in extreme cases destroying the economy’s topology resulting in severe economic recession. This is a completely avoidable situation. The bank should be printing the money that it is loaning, not reducing the amount in the savings accounts.

### v6

@@ VII Solution of the Hamilton-Jacobi-Bellman equation
is quite interesting. Optimal control as done in genAI with Deep Q-Learning (DQN) (Mnih et al., 2015) is based on parametric estimations of a Q-function
@@
Given this theory, we move on to the practical application of it to control the system. This application will use the concepts of Artificial Intelligence (Hastie et al., 2009; Sugiyama, 2015; Goodfellow et al., 2016), as interpreted by Glinsky (2024b, c). First, construct a dataset by either doing an ensemble of simulations of the system or by observing the system. It will be assumed that the system has a small number of dimensions. Most of the systems of interest present themselves in the domain of collective fields $f(x)$ that are elements of a Hilbert space not $q(\tau)$ with a small number of components. How to de-convolute from the domain of the collective motion, the field $f(x)$, to the domain of the individual, $q(\tau)$, (that is from a Hilbert space to $\mathbb{C}^{n}$) using the Heisenberg Scattering Transformation (HST) and a Principal Components Analysis (PCA) will be discussed at the end of this section. This transformation can be done because the collective acts as one, because of the correlation or synchronization specified by the S-matrix $S_{m}$ given in Eq.  91, which are the derivatives of the analytic Hamiltonian $H(\beta)$ of the individual. It is helpful to apply an external force $F_{\text{ext}}(p,q)$ to the system being simulated or observed to sample phase space more efficiently. A good choice would be a dissipation or a random diffusion which samples phase space well, as the system gradually relaxes to the stable equilibriums. It is also good to apply an external force that is constructed to keep the dynamic trajectory in the vicinity of the unstable local maximums, that is stabilizes the unstable equilibriums. The set of variables that should be recorded are $\tau$, in addition to variables that are related to the state $q(\tau)$ and the co-state $p(\tau)$.
@@
Given the dataset, train a neural network with an architecture that matches the structure of the solution to the HJB equation to estimate: (1) the decoding of the $p$ and $q$ coordinates into the $P(p,q)$ and $Q(p,q)$ coordinates that are the solution to the HJB equation as well as the encoding of $P$ and $Q$ to $p(P,Q)$ and $q(P,Q)$, (2) the value function $W(q,P)$ that generates these canonical transformations and is related to the imaginary part of the analytic Hamiltonian as shown in Eq. 83, (3) the mapping of $P$ to the real part of the analytic Hamiltonian $E_{P}(P)$, (4) the frequency $\omega_{Q}(P)$, (5) the policy $\pi(q,P)$ and (6) the analytic advance of $P$ and $Q$ given in Eqns. 79 and 78. Multi Layer Perceptrons (MLPs) (Goodfellow et al., 2016) are used to approximate some of the functions. The derivative functions are calculated by back propagating the MLPs. This architecture is shown in Fig. 9. It is important that Rectified Linear Units (ReLUs) are used as activation functions in the MLPs because the MLPs are approximating analytic functions which are maximally flat but do have a limited number of singularities where the derivative is discontinuous. MLPs with ReLUs are very good at doing this since they are universal piecewise linear approximators with discontinuities in the derivative.
@@
Figure 9: The neural network architecture to estimate the solution of the Hamilton-Jacobi-Bellman equation. It starts with a decoder from $(p,q)$ to $(P,Q)$ generated by the value function $W(q,P)$ related to the imaginary part of the analytic function $H(\beta)$ by Eq. 83, a mapping to the real part of the analytic function $H(\beta)$ given by $E_{P}(P)$, the frequency $\omega_{Q}(P)$, the policy $\pi(q,P)$, the analytic mapping to advance $(P,Q)$, and a encoder from $(P,Q)$ to $(p,q)$. The universal function approximator is a Multi Layer Perceptron (MLP).
@@
Given the function $E_{P}(P)$ estimated in the workflow shown in Fig. 9, $P^{*}$ can be found with a high performance root finder, both stable and unstable equilibriums. The equilibrium policy can then be estimated as $\pi^{*}(q)=\pi(q,P^{*})$ and the equilibrium value as $V^{*}(q)=W(q,P^{*})$. One now has estimated the $\beta^{*}$ which are viewed different ways by different technical disciplines. These are: (1) the ground states of quantum field theory (Weinberg, 2005), (2) the attractive manifolds of nonlinear dynamics (Lichtenberg and Lieberman, 2010), (3) the emergent behaviours and self organizations of complex systems (Gros, 2015), (4) the Taylor relaxed states (Taylor, 1986) and BGK modes (Bernstein et al., 1957) of plasma physics, (5) the poles and branch cuts of control theory and complex analysis (Nehari, 2012), and fundamentally (6) the homology classes of the topology of the dynamic manifold or the geometry of the physics (Frankel, 2012). The equilibrium values $\beta^{*}$ need not be points. They can be manifolds with rich topography, that is algebraic structures, if $n>1$.
@@
There are also unstable equilibriums where $\omega^{\prime}_{Q}<0$, that is local maximums. These unstable equilibriums are also known as saddle points or x-points or semi-attractors that first attract the trajectory to them with the system staying near them for a long time since $\omega_{Q}=0$ (a meta stable state). But, eventually the system will start to move away from the semi-attractor, being repelled away from it. The trajectory will eventually approach and be attracted back to the semi-attractor or might approach and be attracted to another semi-attractor (or several other semi-attractors) before being repelled by the last semi-attractor and returning to the first semi-attractor. An example of a set of two semi-attractors could be a bull and a bear market. These saddle points will need to be stabilized by application of an external conservative force, then cooled by a matched non-conservative thermal force, otherwise the trajectory will evolve into the basin of attraction of a stable attractor and eventually relax to that attractor. While small amounts of energy are needed to move from semi-attractor to semi-attractor, a significant amount of energy is needed to move from one attractor to another attractor or semi-attractor. One needs to climb out of the basin of attraction or potential well, that is climb back up to the mountain pass (saddle point) out of the basin before descending into another basin. When these semi-attractors are stabilized, there is still a local minimization of the economic action to efficiently accomplish the task, but the sub-economy is kept at a global sustainable maximum.
@@
Figure 14: Graph showing the energies $E(P^{*})$, where $\omega_{Q}(P^{*})=0$, as a function of the regularization or dissipation $\nu$ and the strength of the ponderomotive stabilization force $f_{0}$. The thick green line shows the unstable equilibrium or x-point or local maximum at infinity with energy $E_{\text{x}\infty}$. The thick red line shows the stable equilibrium or o-point or Nash Equilibrium at zero with energy $E_{\text{o0}}$. This is the point of severe economic depression. The thin green line shows the local maximum at $P^{*}_{\text{x}}$ with energy $E_{\text{x}}$. The thin red line shows the Nash Equilibrium at $P^{*}_{\text{o}}$ with energy $E_{\text{o}}$. This is the point of economic recession. The thin dashed green lines show the local maximums of the stabilization force with energies $E_{\text{+}}$ and $E_{-}$. The thin dashed red line shows the Nash Equilibrium of the stabilization force with energy $E_{\text{s}}$. This is the point of economic prosperity. The upper limit on viscosity so that the singularities $P^{*}_{\text{x}}$ and $P^{*}_{\text{o}}$ are not destroyed is $\nu<\nu_{\text{cr}}$.
@@
Now to the postponed, but important issue, that the system does not present itself in the low finite dimensional dynamical state $q$ and canonically conjugate momentum $p$ or co-state of the individual, but as an infinite dimensional field $f(x)$ and its canonically conjugate field momentum $\pi(x)$ of the collective (Glinsky, 2024b, c), where $x$ is the base manifold or space with metric $\tau(x)$. Another way of looking at this is that nature presents itself as a convoluted form of the dynamical variables that needs to be de-convoluted or decoded into the dynamical variables. It is very common that $x=t$ and $\tau(t)=t$, but could be as sophisticated as $x=(x,y,z,t)$ and $\tau(x,y,z,t)=\sqrt{t^{2}-(x^{2}+y^{2}+z^{2})/c^{2}}$ (that is relativistic dynamics on space-time). The transformation from the infinite dimensional Hilbert space with coordinates $(\pi,f)$ to the finite dimensional $\mathbb{C}^{n}$ space with coordinates $(p,q)$, has been an ongoing challenge to physics and complex system analysis. It has manifested itself as the renormalization challenge to physicists that was first addressed by Ken Wilson (Wilson, 1971) and has been the subject of two Nobel prizes. Paul Dirac felt that renormalization is “not a logical mathematical process” (Dirac, 1982). There have been recent developments on this subject that have approached renormalization as a logical mathematical process, culminating in the Heisenberg Scattering Transformation (HST). The HST is a closed form, specified Convolutional Neural Network (CNN) or Wavelet Conditional Renormalization Group (WC-RG) (Marchand et al., 2022) which is a transformation from canonical $(\pi,f)$-field phase space to $\mathbb{C}^{n}$ with coordinates $\beta$. When combined with the specification of the analytic function $H(\beta)$ or the equilibrium points (topological homology classes) $\beta^{*}$, it is a Generative Pretrained Transformer (GPT) (Li et al., 2022), and when further deployed as a controller it is a Deep Reinforcement Learning (DRL) (Bertsekas and Tsitsiklis, 1996; Sutton and Barto, 2018; Yoon et al., 2021; Wang et al., 2021; Mnih et al., 2015). The HST will be used as a pre-processor, and the inverse HST (iHST) will be used as a post-processor to the previously discussed analysis, as shown in Fig. 16. A detailed discussion of the relationship of our optimal control methodology to DRL can be found in Sec. X, where particular attention is paid to practical differences between our optimal control methodology and DRL.
@@
Details of the HST can be found in Glinsky and Maupin (Glinsky and Maupin, 2023; Glinsky, 2024a, 2023a, 2023b), along with an example of how it can be combined with an MLP and PCA to form a surrogate model for resistive MagnetoHydroDynamics. The equation for the HST is
@@
What the control method of this section does is move the equilibriums from their uncontrolled natural positions to the desired positions according to the reward function $R(q)$, not $R_{0}(q)$, and stabilizes and cools a more desirable (with more economic activity) saddle point or semi-attractor.
We digress to note the present use of the prime interest rate $\nu_{\text{prime}}$ in economic control. One of the ways that the money supply is traditionally controlled is by a central reserve bank loaning money at a set rate. If this rate is lower than the natural prime interest rate, the money supply will increase as money is borrowed from the central reserve bank. This is what is called quantitative easing and puts inflationary pressure on the economy. The money supply will decrease if the set rate is greater than the prime rate. This reduction in money supply will counter inflation. As will be discussed in Sec. XVI, this is a poor control system since the loaning of money causes a long-term consistent reduction in the money supply due to the coupon payments, countering the original increase in the money supply.
@@ VIII FASER: the LASER of financial systems
The financial FASER works the same way. The financial sub-system needs to be in a meta stable excited state, enjoying the irrational exuberance of a bull market. This meta stable state is likely an unstable saddle point or semi-attractor that takes an infinite amount of time to approach since $\omega_{Q}^{*}=0$. The market is then pumped into an even more excited bull state that is so over priced that it triggers a bear market and the market falls back into its ground (that is) stable state. The financial gain (or profit/loss) is $E_{\text{gain}}$ which is positive for $\Delta E_{p}<\Delta E_{d}$. If the system is already in the ground state $E_{\text{gain}}=-\Delta E_{p}<0$ – a guaranteed money losing proposition pumping energy from the entity doing the “pump and dump” to the financial sub-system.
@@ X Relationship to Deep Reinforcement Learning
Once these surrogates are formed, the equilibrium points $P^{*}$ can be efficiently found in 10’s of seconds. Not only are the stable equilibriums (that is local minimums, o-points, or attractors) found, but the more desirable unstable equilibriums (that is local maximums, x-points, saddle points, or semi-attractors) are also found. These unstable equilibriums can then be stabilized and protected from external forces by the use of Eq. 100 or Eq. 99. The reward function can also be modified as shown in Eq. 88. DRL uses a much slower gradient decent that needs to calculate two time-consuming minimums or maximums for each sample.
@@ XI Quantum field nature of financial systems
The quantum field structure of financial systems comes about from real-time or space-like measurement limitations to those of a statistical nature, so that the Born Rule is applicable. The Born Rule states that if a statistical measurement is made on a system, the measurement must exert a force on the system that will change the system dynamics – the measurement is destructive. Only statistical, that is destructive measurements, can be made on relativistic systems, systems that present themselves in the convoluted field domain of the collective, or systems where non-destructive observations are not available. Practically, this comes about from the time that it takes to communicate, sum, or assimilate the measurements. For instance, financial statements are only issued quarterly or annually, in a retrospective nature. Even if they are reported real-time, they are still only retrospective. The destructive measurement forces are most often non-conservative viscous forces, but can be conservative forces that erect a potential barrier that redirect, not extract energy from, the dynamical trajectory.
@@
The current banking system suffers from an inability to make detailed non-destructive observation of the system. It results in many different money supplies $M$, that are all partial or indirect estimates of the true money supply, at best. The true money supply must be destructively measured as discussed earlier in this section. It is aggravated by the liquidity uncertainty generated by the banks not holding most of the balance of its customers savings accounts in cash. This requires that the depositors treat their deposits statistically, and make statistical measurements. The common statistical measurement of the depositors is requiring the bank to make regular interest payments. In more extreme cases, it can result in the depositors making a very destructive measurement, a bank run where the depositors require the bank to return all their deposits, that results in bank failures, that is bank death, as discussed in Sec. XVI. In both cases, dissipation is added to the economy, degrading the economy’s geometry and performance, and in extreme cases destroying the economy’s topology resulting in severe economic recession. This is a completely avoidable situation. The bank should be printing the money that it is loaning, not reducing the amount in the savings accounts.